A horizontal arm, capable of rotation about a vertical shaft, carries a mass of 2.5 kg, bolted to it by a 1 cm bolt at a distance of 50 cm from the axis of the shaft. The axis of the bolt is vertical. if the ultimate shearing strength of the bolt is 50 MN/ m 2, at what speed will the bolt snap?




1
Expert's answer
2021-05-07T08:13:00-0400

From the formula of circular vertical motion

where T= 50 MN/ m 2

and m = 2.5 Kg , g= 10 m/s"^2" r = 50cm

"T= mg + \frac{mv^2}{r}"


"50= 2.5 \times 10 + \frac{m2.5v^2}{0.5}"


50= "5v^2 + 25"

5"v^2" =25

v= 2.236 m/s